Could you help me solve this problem?
The interest rate on an auto loan in May was 12 3/8%. By September the rate was up to 14 1/4%. How many percentage points did the interest rate increase over the period?
Thanks!

Answer by fractalier(6550)

14 1/4% - 12 3/8% =
14 2/8% - 12 3/8% =
13 10/8% - 12 3/8% =
1 7/8%